# Quantum Frontiers in High Energy Physics

### v1

@@ IV Quantum Machine Learning
HEP experiments generate data of immense complexity and volume, posing significant challenges in data analysis for the objectives of precision measurements and new physics hunt. Machine learning, as a transformative tool, has been extensively studied in HEP data analysis to enhance the overall performances regarding training efficiency and detection accuracy, see a live reviewed in [288]. As facilities like the upcoming HL-LHC and future colliders will produce more than ten times larger datasets, these classical ML methods might become increasingly strained. Conversely, quantum machine learning (QML), leveraging its parallel processing capabilities and quantum properties such as entanglement, may offer the potential to reduce computational resources as well as further improving the overall performances in HEP data analysis. Extensive work in this direction has emerged recently, exploring the potentials of QML in HEP and looking for quantum advantages in the context of speedup in training and improved performance, as reported in [289, 290, 291]. In the following, we provide a concise summary of these studies, which encompass a wide range of data analysis tasks, including object reconstruction, quantum generative models, classification, and anomaly detection, highlighting the perspectives they offer related to quantum advantages, and discuss future directions at the end
@@
In each collision, the produced quarks and gluons immediately form collimated particle sprays known as jets. Jet clustering, which groups final-state particles into jets, is crucial as it retains the information of the originating quark or gluon and forms the basis for studying properties of the Higgs boson, as well as searching for new physics. Several studies [292, 293, 294] have pursued potential quantum advantage in jet clustering, using quantum k-Means/$k_{T}$ algorithm or quantum annealing. By mapping collision events into graphs—with particles as nodes and their angular separations as edges, Ref [295] presents a novel quantum realization of the jet clustering using the Quantum Approximate Optimization Algorithm (QAOA). With up to 30 qubits used on a noiseless quantum simulator, the performance of the QAOA method, based on the angle between the reconstructed jet and the corresponding quark—which ideally should be collimated—was found to be comparable to that of the classical $e^{+}e^{-}k_{t}$ algorithm and superior to that of the classical k-Means, as shown in Fig. 6 from [295], The QAOA methods are further implemented on the Baihua quantum real machine on Quafu [239] using 6 qubits. Mapping jets to graphs also allows jet clustering to be formulated as a quadratic unconstrained binary optimization (QUBO) problem, which quantum algorithms are expected to solve efficiently. A QUBO-based multijet reconstruction has recently been implemented using novel quantum annealing-inspired algorithms (QAIA) [296]. These study highlights the feasibility of quantum computing to revolutionize jet clustering.
Figure 6: Jet clustering performance for 6-particle events on the Baihua quantum hardware is compared to that of quantum simulators, the classical $e^{+}e^{-}k_{t}$ algorithm, and the classical k-Means [295].
@@
Advances have been seen recently in the development of quantum-assisted models for Monte Carlo event generation and for detector simulation. It is shown that the quantum generator architectures in the context of GAN for Monte Carlo event generation [304] are able to learn the kinematic distributions of particles produced at LHC together with the benefits of higher training efficiency in terms of network depth. Complex correlations in the data can also be captured by quantum generative models such as quantum circuit-born machines (QCBM) that are difficult for classical models to learn, as studied in [305]. Detector simulations with assistance from quantum algorithms are mainly focusing on simulating the calorimeter response, where reproducing the images of a reduced sizes of calorimeter responses has been achieved [306]. Prototypes of quantum generative models with continuous variable quantum computing are being suggested with the necessity of further explorations to demonstrated its efficacy [307]. These algorithms on NISQ devices are being tested where optimizations are needed to reduce the effects of quantum noises [308, 309].
@@
Starting from the application of quantum annealing machine learning to optimize the classifiers for distinguishing Higgs events and background events in Ref. [310], exploring the feasibility and advantages of QML methods in the classification tasks has received increasing attentions. Modifications to the quantum annealing machine learning in classification tasks are studied, reducing the performance gap to the its classical counterpart at large training sample sizes [311]. Two different QML methods –Quantum variational classifier and support vector machine with a quantum kernel estimator (QSVM-Kernel method) – have been applied to a flagship analysis (Higgs boson production in association with a top quark pair) at LHC [312, 313], where comparable performance with classical machine learning methods are observed. These QML methods including hybrid quantum-classical graph convolutional network (QGCNN) have also been explored in jet classification [314], particle identifications [315, 316, 317, 318], measurements of physical processes at existing and future lepton colliders [319, 320], where comparable discrimination power to other traditional machine learning models are usually found. A variational quantum searching neighbor algorithm has also been proposed for identifying signals from new physics in the framework of effective field theories [321]. Despite the limitations of quantum computers, such as a limited number of qubits and high quantum error rates, the feasibility of quantum machine learning in classification tasks has begun to be established with these comprehensive studies.
@@
Anomaly detection problem — The absence of new physics discoveries at the LHC has motivated the exploration of model-independent data-analysis methods to efficiently identify anomalous events that could provide hints of new physics. Unsupervised machine learning techniques, also referred to as anomaly detection, are thus being widely studied [288]. Two neural network approaches—autoencoders, which are trained to learn only background distributions, and weakly supervised methods, which are trained to distinguish data from background-only samples—have become major research focuses in anomaly detection. The quantum counterpart of anomaly detection are also emerging along the advances of QML. Quantum anomaly detection based on autoencoder are explored to detect anomalous events at LHC [322, 323, 324, 325], where better performance in the detection accuracy are mostly observed over their corresponding classical methods [322, 323, 324]. Interestingly, it has also been demonstrated that such advantages can be attributed to the inherent properties of the quantum circuits, such as quantum entanglements [323, 324]. Different methods of embedding data into qubits are also being explored to further improve the detection accuracy as well as the representability of the data [326]. Nevertheless, studies that apply quantum anomaly detection in the context of weakly supervised learning—specifically in the four lepton final state at the LHC—have yet to show clear advantages, which might be due to the specific nature of the problem, the limited size of the training dataset, or the particular quantum algorithms used [327].
@@ V Probing quantum nature at colliders
The research scope of HEP has expanded significantly at the high-energy frontier, thanks to systematic measurements and searches across thousands of sensitive phase spaces following the discovery of the Higgs boson in 2012. One intriguing research direction in HEP that has emerged is the study of the quantum nature (QN) of microscopic interactions at high-energy colliders. This allows for an exploration of the fundamental properties of quantum mechanics in the unprecedented high-energy regime, offering a potential tool or even a new philosophical framework that could guide physicists in making future advancements in the search for physics beyond the Standard Model (BSM). The study of quantum mechanics in HEP experiments focuses on fundamental properties such as quantum entanglement and the Bell inequality [328], which are key features of quantum states and some of the most distinctive aspects of quantum mechanics.
@@
In practice, entanglement witnesses that can be easily computed are usually considered, and one widely used example is the concurrence $\mathcal{C}[\rho]$, with its value ranging from zero (for separable, unentangled states) to 1 (for maximally entangled states). For bipartite systems of a pair of qubits ($n=2\times 2$), $\mathcal{C}[\rho]$ is analytically calculable [332, 333]. It can be found using the auxiliary $4\times 4$ matrix [333]
$R=\rho\,(\sigma_{y}\otimes\sigma_{y})\,\rho^{*}\,(\sigma_{y}\otimes\sigma_{y})\,,$ (3)
where $\rho^{*}$ is the complex conjugate of $\rho$. Although non-Hermitian, $R$ possesses non-negative eigenvalues; denoting their square roots as $r_{i}$ with $i=1,2,3,4$, and assuming $r_{1}$ to be the largest, $\mathcal{C}[\rho]$ can be expressed as [333]
@@
For bipartite systems of dimension larger than $2\times 3$, $\mathcal{C}[\rho]$ is not analytically calculable. Fortunately, lower bounds on $\mathcal{C}[\rho]$ for a generic density matrix $\rho$ have been determined and, if non-vanishing, unequivocally signal the presence of entanglement. One of these bounds is computable [334]
$\big{(}\mathcal{C}[\rho]\big{)}^{2}\geq\mathcal{C}_{2}[\rho]\ ,$ (5)
where
@@
$\rho$ $=$ $\frac{1}{9}\left[\mathbb{1}\otimes\mathbb{1}\right]+\sum_{a=1}^{8}f_{a}\left[T^{a}\otimes\mathbb{1}\right]$ (14)
$+$ $\sum_{a=1}^{8}g_{a}\left[\mathbb{1}\otimes T^{a}\right]+\sum_{a,b=1}^{8}h_{ab}\left[T^{a}\otimes T^{b}\right]\,,$
@@
Phenomenological studies and expected sensitivities — The procedure to determine the density matrix of a quantum state by performing measurements over an ensemble of events is known as quantum tomography [339, 340, 341]. This method has been widely used in collider studies, and a complete formalism for the quantum tomography of the density matrix $\rho$ in general scattering processes has been developed in Ref. [342, 343]. The coefficients in the parametrization of $\rho$ can, in principle, be determined by averaging over the angular distribution of the final state in dedicated reference frames. Certain aspects unique to high-energy experiments need to be highlighted. In colliders, analyses allow each event to be reconstructed in different frames. Collecting the ensemble of events may not result in a quantum state, if the averaging is done in event-dependent frames. The object constructed in this manner has been labeled a “fictitious state” [344], which loses many characteristic properties of quantum states. Given the frame dependence on a per-event basis in collider experiments, there is an opportunity to determine optimal frame choices for maximizing spin correlations, entanglement, and Bell inequality violations [345, 346]. While the decays of particles are mostly considered for quantum tomography, a novel approach, referred to as the kinematic approach was recently proposed [347]. This approach relies solely on the kinematics of particles to perform quantum tomography. By leveraging the simplicity of scattering kinematics, this approach promises optimal statistical performance and offers the opportunity to establish Bell non-locality within existing data. Building on these developments, along with those summarized in Table 3, promising sensitivities can be achieved using existing or foreseen data that will be accumulated. Below, we detail one example of measuring quantum entanglement at ATLAS and CMS.
@@
$\tau$-lepton pair production at the LHC [355] and SuperKEKB [356] Expected to observe both QE and BI violation at $5~{}\sigma$ with exiting Belle II data [355]
@@
Diboson ($WW,WZ,ZZ$) production at LHC [343, 361, 362] Observation of QE might be possible BI violation not expected
@@
The second topic is the measurement of final-state particle polarization at colliders. Measuring boson polarization has a long history since the LEP and Tevatron experiments, where the focus was on the abundantly produced single vector boson process ($V=W,Z$) or a vector boson in association with jets ($V+$ jets). Polarization measurements in these channels significantly improve the understanding of electroweak and QCD theories, and in particular, improve the precision in determining fundamental parameters such as the weak mixing angle, as well as constraining parton distribution functions and higher-order QCD calculations. The production of two or more bosons (multi-bosons) is key to probing the gauge structure of the SM and searching for new physics, but this is not well measured due to insufficient statistics and limited collision energy. In the LHC era, the abundant high-energy collision data provides a unique opportunity for substantial progress in this area, and in fact, adding the polarization dimension to measurements helps to sharpen the physics sensitivity to the next level. This can be easily understood from various physics discussions, for example: heavy resonances in many new physics models preferentially decay into two longitudinally polarized vector bosons ($V_{L}V_{L}$); the longitudinally polarized vector boson is directly related to the Goldstone boson equivalence theorem, which exhibits unique features enabling the probe of the relevant physics in greater depth.
@@ Acknowledgements.
We would like to thank Tao Han, Henry Lamm, Hideki Okawa and Yue Zhao for their insightful feedback and valuable comments on our manuscript. Y. Fang is supported by the Institute of High Energy Physics, Chinese Academy of Sciences under the innovative project on sciences and technologies with Grant No. E3545BU210. Y.-Y. Li is supported by the NSF of China through Grant No. 12305107, No. 12247103. J. Shu is supported by Peking University under startup Grant No. 7101302974 and the NSF of China under Grants No. 12025507, No.12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences under Grants No. ZDBS-LY-7003. Y. Wu and L. Xu are supported by the NSF of China International Cooperation Grant No. 12061141004, NSF of China Grant No. 12275265, and MOST National Key Technologies R&D Program No. 2023YFA1605703. H. Xing is supported by the NSF of China with Grant No. 12035007 and Guangdong Major Project of Basic and Applied Basic Research with No. 22020B0301030008, 2022A1515010683. C. Zhou is supported by the Fundamental Research Funds for the Central Universities, Peking University.

### v2

@@ IV Quantum Machine Learning
HEP experiments generate data of immense complexity and volume, posing significant challenges in data analysis for the objectives of precision measurements and new physics hunt. Machine learning, as a transformative tool, has been extensively studied in HEP data analysis to enhance the overall performances regarding training efficiency and detection accuracy, see a live reviewed in [312]. As facilities like the upcoming HL-LHC and future colliders will produce more than ten times larger datasets, these classical ML methods might become increasingly strained. Conversely, quantum machine learning (QML), leveraging its parallel processing capabilities and quantum properties such as entanglement, may offer the potential to reduce computational resources as well as further improving the overall performances in HEP data analysis. Extensive work in this direction has emerged recently, exploring the potentials of QML in HEP and looking for quantum advantages in the context of speedup in training and improved performance, as reported in [313, 314, 315]. In the following, we provide a concise summary of these studies, which encompass a wide range of data analysis tasks, including object reconstruction, quantum generative models, classification, and anomaly detection, highlighting the perspectives they offer related to quantum advantages, and discuss future directions at the end.
@@
In each collision, the produced quarks and gluons immediately form collimated particle sprays known as jets. Jet clustering, which groups final-state particles into jets, is crucial as it retains the information of the originating quark or gluon and forms the basis for studying properties of the Higgs boson, as well as searching for new physics. Several studies [316, 317, 318] have pursued potential quantum advantage in jet clustering, using quantum k-Means/$k_{T}$ algorithm or quantum annealing. By mapping collision events into graphs—with particles as nodes and their angular separations as edges, Ref [319] presents a novel quantum realization of the jet clustering using the Quantum Approximate Optimization Algorithm (QAOA). With up to 30 qubits used on a noiseless quantum simulator, the performance of the QAOA method, based on the angle between the reconstructed jet and the corresponding quark—which ideally should be collimated—was found to be comparable to that of the classical $e^{+}e^{-}k_{t}$ algorithm and superior to that of the classical k-Means, as shown in Fig. 6, The QAOA methods are further implemented on the Baihua quantum real machine on Quafu [262] using 6 qubits. Mapping jets to graphs also allows jet clustering to be formulated as a quadratic unconstrained binary optimization (QUBO) problem, which quantum algorithms are expected to solve efficiently. A QUBO-based multijet reconstruction has recently been implemented using novel quantum annealing-inspired algorithms (QAIA) [320]. These study highlights the feasibility of quantum computing to revolutionize jet clustering.
Figure 6: Figure taken from  [319]:Jet clustering performance for 6-particle events on the Baihua quantum hardware is compared to that of quantum simulators, the classical $e^{+}e^{-}k_{t}$ algorithm, and the classical k-Means.
@@
Advances have been seen recently in the development of quantum-assisted models for Monte Carlo event generation and for detector simulation. It is shown that the quantum generator architectures in the context of GAN for Monte Carlo event generation [328] are able to learn the kinematic distributions of particles produced at LHC together with the benefits of higher training efficiency in terms of network depth. Complex correlations in the data can also be captured by quantum generative models such as quantum circuit-born machines (QCBM) that are difficult for classical models to learn, as studied in [329]. Detector simulations with assistance from quantum algorithms are mainly focusing on simulating the calorimeter response, where reproducing the images of a reduced sizes of calorimeter responses has been achieved [330]. Prototypes of quantum generative models with continuous variable quantum computing are being suggested with the necessity of further explorations to demonstrated its efficacy [331]. These algorithms on NISQ devices are being tested where optimizations are needed to reduce the effects of quantum noises [332, 333, 334].
@@
Starting from the application of quantum annealing machine learning to optimize the classifiers for distinguishing Higgs events and background events in Ref. [335], exploring the feasibility and advantages of QML methods in the classification tasks has received increasing attentions. Modifications to the quantum annealing machine learning in classification tasks are studied, reducing the performance gap to the classical counterpart at large training sample sizes [336]. Two different QML methods –Quantum variational classifier and support vector machine with a quantum kernel estimator (QSVM-Kernel method) – have been applied to a flagship analysis (Higgs boson production in association with a top quark pair) at LHC [337, 338], where comparable performance with classical machine learning methods are observed. These QML methods including hybrid quantum-classical graph convolutional network (QGCNN) have also been explored in jet classification [339], particle identifications [340, 341, 342, 343], measurements of physical processes at existing and future lepton colliders [344, 345], where comparable discrimination power to other traditional machine learning models are usually found. A variational quantum searching neighbor algorithm has also been proposed for identifying signals from new physics in the framework of effective field theories [346]. Despite the limitations of quantum computers, such as a limited number of qubits and high quantum error rates, the feasibility of quantum machine learning in classification tasks has begun to be established with these comprehensive studies.
@@
Anomaly detection problem — The absence of new physics discoveries at the LHC has motivated the exploration of model-independent data-analysis methods to efficiently identify anomalous events that could provide hints of new physics. Unsupervised machine learning techniques, also referred to as anomaly detection, are thus being widely studied [312]. Two neural network approaches—autoencoders, which are trained to learn only background distributions, and weakly supervised methods, which are trained to distinguish data from background-only samples—have become major research focuses in anomaly detection. The quantum counterpart of anomaly detection are also emerging along the advances of QML. Quantum anomaly detection based on autoencoder are explored to detect anomalous events at LHC [347, 348, 349, 350], where better performance in the detection accuracy are mostly observed over their corresponding classical methods. Interestingly, it has also been demonstrated that such advantages can be attributed to the inherent properties of the quantum circuits, such as quantum entanglements [348, 349]. Different methods of embedding data into qubits are also being explored to further improve the detection accuracy as well as the representability of the data [351]. Improvements have also been observed recently in anomaly detection using quantum similarity learning, surpassing its counterpart [352]. Nevertheless, studies that apply quantum anomaly detection in the context of weakly supervised learning—specifically in the four lepton final state at the LHC—have yet to show clear advantages, which might be due to the specific nature of the problem, the limited size of the training dataset, or the particular quantum algorithms used [353].
@@ V Probing quantum nature at colliders
The research scope of HEP has expanded significantly at the high-energy frontier, thanks to systematic measurements and searches across thousands of sensitive phase spaces following the discovery of the Higgs boson in 2012. One intriguing research direction in HEP that has emerged is the study of the quantum nature (QN) of microscopic interactions at high-energy colliders [354]. This allows for an exploration of the fundamental properties of quantum mechanics in the unprecedented high-energy regime, offering a potential tool or even a new philosophical framework that could guide physicists in making future advancements in BSM searches. The study of quantum mechanics in HEP experiments focuses on fundamental properties such as quantum entanglement and the Bell inequality [355], which are key features of quantum states and some of the most distinctive aspects of quantum mechanics.
@@
In practice, entanglement witnesses that can be easily computed are usually considered, and one widely used example is the concurrence $\mathcal{C}[\rho]$, with its value ranging from zero (for separable, unentangled states) to 1 (for maximally entangled states). For bipartite systems of a pair of qubits ($n=2\times 2$), $\mathcal{C}[\rho]$ is analytically calculable [359, 360]. It can be found using the auxiliary $4\times 4$ matrix $R=\rho\,(\sigma_{y}\otimes\sigma_{y})\,\rho^{*}\,(\sigma_{y}\otimes\sigma_{y})$ [360] with $\rho^{*}$ being the complex conjugate of $\rho$. Although non-Hermitian, $R$ possesses non-negative eigenvalues; denoting their square roots as $r_{i}$ with $i=1,2,3,4$, and assuming $r_{1}$ to be the largest, $\mathcal{C}[\rho]$ can be expressed as [360]
@@
For bipartite systems of dimension larger than $2\times 3$, $\mathcal{C}[\rho]$ is not analytically calculable. Fortunately, lower bounds on $\mathcal{C}[\rho]$ for a generic density matrix $\rho$ have been determined and, if non-vanishing, unequivocally signal the presence of entanglement. One computable bound among these is $\big{(}\mathcal{C}[\rho]\big{)}^{2}\geq\mathcal{C}_{2}[\rho]$ [361] where
@@
$\rho$ $=$ $\frac{1}{9}\left[\mathbb{I}\otimes\mathbb{I}\right]+\sum_{a=1}^{8}f_{a}\left[T^{a}\otimes\mathbb{1}\right]$ (12)
$+$ $\sum_{a=1}^{8}g_{a}\left[\mathbb{I}\otimes T^{a}\right]+\sum_{a,b=1}^{8}h_{ab}\left[T^{a}\otimes T^{b}\right]\,,$
@@
Phenomenological studies and expected sensitivities — The procedure to determine the density matrix of a quantum state by performing measurements over an ensemble of events is known as quantum tomography [366, 367, 368]. This method has been widely used in collider studies, and a complete formalism for the quantum tomography of the density matrix $\rho$ in general scattering processes has been developed in Ref. [369, 370]. The coefficients in the parametrization of $\rho$ can, in principle, be determined by averaging over the angular distribution of the final state in dedicated reference frames. Certain aspects unique to high-energy experiments need to be highlighted. In colliders, analyses allow each event to be reconstructed in different frames. Collecting the ensemble of events may not result in a quantum state, if the averaging is done in event-dependent frames. The object constructed in this manner has been labeled a “fictitious state” [371], which loses many characteristic properties of quantum states. Given the frame dependence on a per-event basis in collider experiments, there is an opportunity to determine optimal frame choices for maximizing spin correlations, entanglement, and Bell inequality violations [372, 373]. While the decays of particles are mostly considered for quantum tomography, a novel approach, referred to as the kinematic approach was recently proposed [374]. This approach relies solely on the kinematics of particles to perform quantum tomography. By leveraging the simplicity of scattering kinematics, this approach promises optimal statistical performance and offers the opportunity to establish Bell non-locality within existing data. Effects of higher-order corrections from QCD and EW in modifying the helicity structure in the production channel are also being scrutinized in [375]. Building on these developments, along with those summarized in Table 3, promising sensitivities can be achieved using existing or foreseen data that will be accumulated. Below, we detail one example of measuring quantum entanglement at ATLAS and CMS.
@@
$\tau$-lepton pair production at the LHC [383], SuperKEKB [384] and FCC-ee [385] Expected to observe both QE and BI violation at $5~{}\sigma$ with exiting Belle II data [383]
@@
B meson decays at LHCb [389, 390] QE expected to be observed at $>5\sigma$ with present LHCb data Expected to test BI violation at $>5\sigma$ with present LHCb data
@@
Diboson ($WW,WZ,ZZ$) production at LHC [370, 392, 393, 394] Observation of QE might be possible BI violation not expected
@@
The second topic is the measurement of final-state particle polarization at colliders. Measuring boson polarization has a long history since the LEP and Tevatron experiments, where the focus was on the abundantly produced single vector boson process ($V=W,Z$) or a vector boson in association with jets ($V+$ jets). Polarization measurements in these channels significantly improve the understanding of electroweak and QCD theories, and in particular, improve the precision in determining fundamental parameters such as the weak mixing angle, as well as constraining parton distribution functions and higher-order QCD calculations. The production of two or more bosons (multi-bosons) is key to probing the gauge structure of the SM and searching for new physics, but this is not well measured due to insufficient statistics and limited collision energy. In the LHC era, the abundant high-energy collision data provides a unique opportunity for substantial progress in this area, and in fact, adding the polarization dimension to measurements helps to sharpen the physics sensitivity to the next level [406]. This can be easily understood from various physics discussions, for example: heavy resonances in many new physics models preferentially decay into two longitudinally polarized vector bosons ($V_{L}V_{L}$); the longitudinally polarized vector boson is directly related to the Goldstone boson equivalence theorem, which exhibits unique features enabling the probe of the relevant physics in greater depth.
@@ Acknowledgements.
We would like to thank Tao Han, Henry Lamm, Hideki Okawa and Yue Zhao for their insightful feedback and valuable comments on our manuscript. Y. Fang is supported by NSFC Basic Science Centre Program for “Joint Research on High Energy Frontier Particle Physics” (Grant No. 12188102); and is supported by the Institute of High Energy Physics, Chinese Academy of Sciences under the innovative project on sciences and technologies with Grant No. E3545BU210. Y.-Y. Li is supported by the NSF of China through Grant No. 12305107, No. 12247103. J. Shu is supported by Peking University under startup Grant No. 7101302974 and the NSF of China under Grants No. 12025507, No.12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences under Grants No. ZDBS-LY-7003. Y. Wu and L. Xu are supported by the NSF of China International Cooperation Grant No. 12061141004, NSF of China Grant No. 12275265, and MOST National Key Technologies R&D Program No. 2023YFA1605703. H. Xing is supported by the NSF of China with Grant No. 12035007 and Guangdong Major Project of Basic and Applied Basic Research with No. 22020B0301030008, 2022A1515010683. C. Zhou is supported by the Fundamental Research Funds for the Central Universities, Peking University.